OpenStudy (anonymous):

solve for t: 4t^2-15t=0

OpenStudy (venomblast):

t=15/4

OpenStudy (anonymous):

Can you display your work? I know factoring comes into play. Thanks a bunch!

OpenStudy (anonymous):

t(4t-15)=0 4t-15=0 4t=15 t=15/4

OpenStudy (anonymous):

So you can just divide 0 by t and it gets rid of it in the first step?

OpenStudy (venomblast):

no. that the zero property. you dont divide by zero or your answer would be undefine

OpenStudy (anonymous):

right, how does the t in the first step dissapear?

OpenStudy (anonymous):

@Venomblast let me tell you... if you divide any number by zero then it will result infinity like 1/0 ... but if you divide 0 by any number you'll get zero as answer... @Topkart33 yes t is divided by zero which results in zero...

OpenStudy (venomblast):

actually. 1/0 = undefine and 0/1 = 0

OpenStudy (venomblast):

@topkart33 the t disapear by taking out a t. so t(4t-15).